Television screens are measured diagonally. If the screens are square, what is the difference in area between a 21-inch screen and a 24-inch screen?
Answer by nerdybill(7384)

Television screens are measured diagonally. If the screens are square, what is the difference in area between a 21-inch screen and a 24-inch screen?
.
21 inch screen: (applying Pythagorean)
x^2 + x^2 = 21^2
2x^2 = 441
x^2 = 220.5 square inches
.
24 inch screen: (applying Pythagorean)
x^2 + x^2 = 24^2
2x^2 = 576
x^2 = 288 square inches
.
Difference in area:
288-220.5 = 67.5 square inches